# Is Deep Learning finally better than Decision Trees on Tabular Data?

### v2

@@ 3.2 Experimental Setup
Metrics. Lastly, we report the model’s performance as the average Area Under the Receiver Operating Characteristic (ROC-AUC) across 10 outer test folds. Given the prevalence of imbalanced datasets in the OpenMLCC18 benchmark, we employ ROC-AUC as our primary metric. This measure offers a more reliable assessment of model performance across varied class distributions, as it is less influenced by the imbalance in a dataset.
In our study, we adhered to the official hyperparameter search spaces from the respective papers for tuning every method. For a detailed description of the hyperparameter search spaces of all other methods included in our analysis, we refer the reader to Appendix A.
Code: For reproducibility, our code is available at: https://anonymous.4open.science/r/TabularStudy-0EE2.
@@ 22       Train final model $M(\lambda^{*}_{k})$ on $D_{\text{train}}^{k}$;
4 Baselines
In our experiments, we compare a range of methods categorized into three distinct groups: Classical Machine Learning Classifiers (§ 4.1), Deep Learning Methods (§ 4.2), and AutoML frameworks (§ 4.3) as shown in Figure 1.
4.1 Classical Machine Learning Classifiers
Gradient Boosted Decision Trees.  First, we consider XGBoost (Chen & Guestrin, 2016), a well-established GBDT library that uses asymmetric trees. The library does not natively handle categorical features, which is why we apply one-hot encoding. Moreover, we consider CatBoost, a well-known library for GBDT that employs oblivious trees as weak learners and natively handles categorical features with various strategies.
4.2 Deep Learning Methods
Dataset-Specific Neural Networks.  Recent works have shown that MLPs featuring residual connections outperform plain MLPs and make for very strong competitors to state-of-the-art architectures (Kadra et al., 2021; Gorishniy et al., 2021). As such, in our study, we include the ResNet implementation provided in Gorishniy et al. (2021). Furthermore, research indicates that incorporating numerical embeddings (Gorishniy et al., 2022), such as PLR, into neural networks significantly improves performance, motivating us to include an MLP with PLR embeddings in our study Throughout this paper, any mention of MLP refers specifically to an MLP with PLR embeddings. Additionally, RealMLP (Holzmüller et al., 2024) not only leverages numerical embeddings but also employs techniques like robust scaling, smooth clipping, and a diagonal weight layer, enabling it to surpass other neural network models. Consequently, we include RealMLP in our analysis as well.
Additionally, we consider several transformer-based architectures designed specifically for tabular data. TabNet (Arik & Pfister, 2021) employs sequential attention to selectively utilize the most pertinent features at each decision step. For the implementation of TabNet, we use a well-maintained public implementation [^2] https://github.com/dreamquark-ai/tabnet. SAINT (Somepalli et al., 2021) introduces a hybrid deep learning approach tailored for tabular data challenges. SAINT applies attention mechanisms across both rows and columns and integrates an advanced embedding technique. Lastly, FT-Transformer Gorishniy et al. (2021) is an adaptation of the Transformer architecture for tabular data. It transforms categorical and numerical features into embeddings, which are then processed through a series of Transformer layers.
Foundation Models for Tabular Classification.  These models can be further divided into two categories based on their learning strategies: In-Context Learning and Fine-Tuning. The former eliminates the need for per-dataset finetuning, whereas the latter requires models to undergo finetuning specific to each dataset.
For in-context learning, we consider TabPFN, a meta-learned transformer architecture.
Among fine-tuned models, XTab proposes a cross-table pretraining approach that can work across multiple tables with different column types and structures. The approach utilizes independent featurizers for individual columns and federated learning to train a shared transformer component, allowing better generalization. Next, TP-BERTa is a variant of the BERT language model being specifically adapted for tabular prediction. It introduces a relative magnitude tokenization to transform continuous numerical values into discrete high-dimensional tokens. Its learning procedure relies on an intra-feature attention module that learns relationships between feature values and their corresponding feature names, effectively bridging the gap between numerical data and language-based feature representation. Lastly, CARTE utilizes a graph representation of tabular data to process tables with differing structures. It applies an open vocabulary setting and contextualizes the relationship between table entries and their corresponding columns by a graph-attentional network. CARTE is pre-trained on a large knowledge base enhancing generalization.
Since all the fine-tuned models were pretrained on real-world datasets, we ensure that no datasets overlapping with the OpenMLCC18 benchmark are included in the evaluation. For all baselines, we use their official implementations. We refer the readers to Appendix A for more details.
4.3 AutoML Frameworks
Due to the large number of AutoML frameworks available in the community (Feurer et al., 2015; Erickson et al., 2020; LeDell & Poirier, 2020; Feurer et al., 2022), it was infeasible to include all of them in our experimental study. Therefore, we selected AutoGluon (Erickson et al., 2020), which is regarded as one of the best AutoML frameworks according to the AutoML Benchmark study (Gijsbers et al., 2024). Unlike other AutoML systems, AutoGluon does not recommend performing hyperparameter optimization, instead, it relies on stacking and ensembling. In our study, we include two versions of AutoGluon: one where we perform hyperparameter optimization for all models, and a version where we follow the original author’s recommendations by setting presets="best_quality".
5 Experiments and Results
In this study, we aim to address several key research questions related to the performance of machine learning techniques and various deep learning model families on tabular data classification:
R1: Do DL models outperform gradient boosting methods in tabular data classification? (§ 5.1)
   $\rightarrow$ GBDTs show robust performance while TabPFN is competitive on small datasets.
R2: Do meta-learned NNs outperform data-specific NNs in tabular data classification? (§ 5.1)
   $\rightarrow$ TabPFN is superior on small datasets. Dataset-specific sota models outperform foundation models.
R3: Do in-context models or fine-tuned models perform better? (§ 5.2)
   $\rightarrow$ TabPFN wins on small datasets. XTab as a fine-tuned model yields the most robust performance.
R4: Do DL models outperform AutoML libraries? (§ 5.3)
   $\rightarrow$ AutoML is superior to DL models on a broad range of datasets.
R5: What is the influence of hyperparameter optimization on the output quality? (§ 5.4)
   $\rightarrow$ HPO shows significant improvements besides for models reaching computational limits.
R6: How do dataset characteristics relate to the performance of different model families? (§ 5.5)
   $\rightarrow$ No significant relations of meta features to performance across different model families.
R7: What is the cost vs. efficiency relation of various model families? (§ 5.6)
   $\rightarrow$ GBDTs show best relations w.r.t. inference time while AutoML is competitive w.r.t. total time.
5.1 (R1+R2) Quality metrics
To address our first two research questions, we compare the performance of different families of methods by ranking each method on every dataset and analyzing the distribution of these ranks. Figure 2 presents a comparison between Deep Learning methods and classical ML methods. The results indicate that Classical ML methods - namely CatBoost and XGBoost - demonstrate robust and consistent performance across datasets, with CatBoost achieving a median rank of 2 and XGBoost a median rank of 2.5. Among the Deep Learning methods, TabPFN exhibits the best performance with a median rank of 2.
Figure 2: Distribution of ranks for the Deep Learning ($10$ methods) and Classical ML ($2$ methods) classifier families. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
Figure 3: Distribution of ranks for the Foundation Models ($4$ methods) and Dataset-Specific ($4$ methods) classifier families. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
To address the second research question R2, we analyzed the distribution of ranks between the two subfamilies within the Deep Learning category: Foundation Models and Dataset-Specific Neural Networks. Figure 3 illustrates that, overall, dataset-specific neural networks outperform foundation models, with the notable exception of TabPFN, which achieved a median rank of 2 across its 17 evaluated datasets. Within the dataset-specific family, RealMLP demonstrated the best performance, attaining a median rank of 3 across all 68 datasets. This is followed by MLP with PLR embeddings and ResNet, both with a median rank of 3, however, MLP exhibited a narrower interquartile range, indicating a more consistent performance. Among the foundation models, after TabPFN, XTab shows the next best performance with a median rank of 3, followed by CARTE, and finally, TP-BERTa, which displays the lowest performance within this group.
To compare foundation models with dataset-specific NNs in the small data regime, we utilized the autorank package Herbold (2020), performing a Friedman test followed by a Nemenyi post-hoc test at a significance level of $0.05$. This statistical analysis allows for generating a critical difference (CD) diagram shown in the left plot of Figure 4.
Figure 4: Statistical analysis in the small data domain (number of instances $\leq 1000$). Left: Critical Difference (CD) Diagram of dataset-specific neural networks against foundation models. Right: Critical Difference (CD) Diagram of Deep Learning models against GBDT models
Due to the limited number of datasets shared among the methods, TP-BERTa was excluded from this comparison. The black bars connecting methods indicate that there is no statistically significant difference in performance. While TabPFN outperforms RealMLP, FT-Transformer, MLP, ResNet, SAINT and XTab these results are not statistically significant. The CD diagram illustrates that in the small data domain, i.e., number of instances $\leq$ 1,000, TabPFN outperforms other methods, demonstrating superior performance. We also compare the performance of deep learning models with classical machine learning models in the small data domain. Right plot of Figure 4 presents a critical difference diagram, which indicates that TabPFN achieves the highest rank overall. A comprehensive presentation of the raw results for all methods, both after hyperparameter optimization (HPO) and with default hyperparameter configurations, is provided in Appendix B.
5.2 (R3) In-context learning vs. Fine-tuning
To further investigate the family of foundation models whether fine-tuning or in-context learning models yield better performance, we conducted an analysis similar to our previous research questions. We employ boxplots to display the distribution of ranks and use critical difference (CD) diagrams to evaluate the statistical significance of the results.
Figure 5 illustrates that TabPFN, categorized under in-context learning methods, achieved a median rank of 1 with no interquartile range, indicating highly consistent performance across its 17 evaluated datasets. Among the fine-tuning methods, XTab showed the best performance with a median rank of 1 but exhibited a larger interquartile range, followed by CARTE and TP-BERTa.
For a statistical comparison, we present a CD diagram in Figure 6, from which TP-BERTa is excluded due to the limited number of common datasets among the methods. The CD diagram reveals that TabPFN outperforms both XTab and CARTE, highlighting its superiority within the Foundation Models category.
Figure 5: Distribution of ranks for the Fine-tuning (3 models) and In-context learning (1 model) classifier families. The boxplots illustrate the rank spread, with medians represented by red lines and whiskers showing the range.
Figure 6: Comparative analysis of Fine-tuning models against In-context learning models in the small data domain (number of instances $\leq 1000$). The horizontal bar indicates the absence of statistically significant differences.
5.3 (R4) Deep Learning models vs. AutoML Libraries
Figure 7: Distribution of ranks for the Deep Learning Models (10 methods) and AutoML (1 method) classifier families. Left: AutoGluon with hyperparameter optimization (HPO). Right: AutoGluon in its recommended configuration. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
Figure 7 presents boxplots showing the distribution of ranks for all Deep Learning methods compared to AutoGluon. The left-hand side illustrates the comparison with the HPO version of AutoGluon. Except for TabPFN, which achieves a median rank of 2, while RealMLP follows closely with a strong median rank of 3, all other methods are outperformed by AutoGluon, which has a median rank of 3, with its interquartile range extending down to rank 1 across its 68 evaluated datasets. The right-hand side displays the recommended version of AutoGluon, which exhibits even better performance with a median rank of 1 and a very narrow interquartile range, indicating robust performance.
Consistent with our previous analyses, we also present critical difference (CD) diagrams for this comparison. Figure 8 shows, on the left side, the CD diagram of AutoGluon with HPO and, on the right side, AutoGluon with its recommended configuration, both compared against the Deep Learning methods across all datasets. The left CD diagram indicates that while AutoGluon attains the highest rank, the differences are not statistically significant when compared to SAINT, FT-Transformer, ResNet, and XTab; however, they are statistically significant when compared to CARTE and TabNet. The right CD diagram tells a different story: AutoGluon with its recommended settings is superior to every method in the comparison with a statistically significant result.
Furthermore, we conduct the same analysis in the small data domain, where the number of instances is $\leq 1000$, to include TabPFN in the analysis. Figure 9 presents these results. color=chris]do we have a reason why auto gluon is that bad in the small data regime? In the left diagram, TabPFN outperforms all other methods, including the HPO version of AutoGluon. However, in the right diagram, AutoGluon with its recommended settings outperforms every method, including TabPFN, nevertheless, the difference is not statistically significant.
5.4 (R5) Analysis of HPO
In our analysis of hyperparameter optimization (HPO) versus default configurations across various machine learning methods, we observed that HPO generally led to improved performance. The analysis is depicted in Figure 10. This improvement is reflected in the average rank reductions for most methods when HPO was applied. For example, XGBoost’s average rank improved significantly from 1.94 in its default configuration to 1.06 with HPO, and XTab showed a similar enhancement, moving from a rank of 1.96 down to 1.04. These findings are visually represented in the accompanying plot, which illustrates the performance gains achieved through HPO. An exception to the general trend was observed with TP-BERTa, where the default configuration slightly outperformed the HPO version (average ranks of 1.47 and 1.53, respectively). This anomaly can be attributed to the computational demands of TP-BERTa. Due to its large model size, TP-BERTa was unable to complete the full 100 hyperparameter tuning trials within the allotted 23-hour time frame, often finishing only a few trials. Consequently, the HPO process may have converged to a suboptimal configuration that did not surpass the performance of the default settings.
(a) CatBoost
(b) XGBoost
(c) ResNet
(d) FT-Transformer
(e) SAINT
(f) TabNet
(g) XTab
(h) CARTE
(i) TP-BERTa
(j) MLP
(k) RealMLP
Figure 11: Hyperparameter importance for various methods
Figure 11 illustrates the importance of the individual hyperparameters tuned for every method. We calculate hyperparameter importance using the fANOVA (Hutter et al., 2014) implementation in Optuna (Akiba et al., 2019). According to our analysis, the most important hyperparameter for CatBoost is the learning rate, while for XGBoost, it is the subsample ratio of the training instances. For XTab, the learning rate is also the most important hyperparameter, closely followed by the light_finetune hyperparameter, which is a categorical parameter taking values True or False. When light_finetune is True, we fine-tune XTab for only 3 epochs; when it is False, we use the same range of epochs as for the other methods (10 to 500). Similarly, for the MLP with PLR embeddings, the learning rate proves to be the most influential hyperparameter, whereas for RealMLP, the number of units in the hidden layers. For the remaining dataset-specific neural networks in the deep learning family, as well as for CARTE, the number of training epochs is the most important hyperparameter, indicating that training duration plays a critical role in their performance.
5.5 (R6) Influence of meta-feature characteristics on the predictive performance
5.6 (R7) Cost vs. efficiency relation
To address the final research question, to see what is the cost vs. efficiency relation of various model families, we plot the intra-search space normalized Average Distance to the Maximum (ADTM) Wistuba et al. (2016) in Figure 14, illustrating how quickly each method converges to its best solution during the HPO process.
The plot shows that XGBoost is the fastest, reaching nearly optimal performance within just 5 hours. ResNet and MLP also demonstrate notable speed, followed closely by CatBoost.
Figure 14: Intra-search space normalized average distance to the maximum over the cumulative training time (seconds).
Overall, the gradient boosting methods (GBDT) converge faster than the deep learning models. XTab, which shares the same transformer architecture as FT-Transformer, exhibits quicker convergence, likely due to its static architecture, while the FT-Transformer’s architectural components were also tuned. On the other hand, TP-BERTa is the slowest to converge, likely due to the high computational demands of its BERT-like architecture.
In Figure 15, we show the performance profiles of the models considered. We first normalize the performance values and the logarithmic time values. Let $P_{i}^{(j)}$ and $T_{i}^{(j)}$ be the performance of algorithm $i$ on dataset $j$, resp., the measured time. Let $m_{\ast}^{(j)}=\max_{i}~{}P_{i}^{(j)}$ be the model yielding the best performance and $t_{\dagger}^{(j)}=\max_{i}~{}T_{i}^{(j)}$ denoting the greatest running time on dataset $j$, we compute the performance gap $\operatorname{gap}_{i}^{(j)}=(m_{\ast}^{(j)}-P_{i}(j))/m_{\ast}^{(j)}$ and the temporal gain $\operatorname{tgain}_{i}^{(j)}=(t_{\dagger}^{(j)}-T_{i}^{(j)})/t_{\dagger}^{(j)}$ being achieved for each algorithm $i$ on a dataset $j$. We define a Performance-Time Ratio $\operatorname{ptr}_{i}^{(j)}=\operatorname{gap}_{i}^{(j)}/\operatorname{tgain}_{i}^{(j)}$ , where $\operatorname{ptr}_{i}^{(j)}>1$ values indicate that the performance gain outweighs the time cost, $\operatorname{ptr}_{i}^{(j)}=1$ yields a balanced view, whereas $\operatorname{ptr}_{i}^{(j)}<1$ indicate that the time cost outweighs the performance gain. We further normalize the $\operatorname{ptr}_{i}^{(j)}$ to be in the range of $[0,1]$ such that values closer to 1 (0) indicate a better (poorer) trade-off in terms of performance gain relative to time cost. To determine the cumulative proportion for a given threshold $\tau$, we count how many normalized $\operatorname{ptr}{{}^{\prime}}_{i}^{(j)}=\operatorname{ptr}_{i}^{(j)}/(1+\operatorname{ptr}_{i}^{(j)})$ are less than or equal to $\tau$ for each algorithm $i$. Hence, the closer a line reaches 1.0 for smaller values of $\tau$, the more often an algorithm is close to having the best performance-time-cost, i.e., lines that rise more steeply indicate better overall ratios on the datasets considered. A red star in the upper left of the illustrations indicates the best value.
On the left, the performance profiles are shown w.r.t. the measured inference time. The evaluation shows that both GBDT models yield the best performance-time ratio, followed by the dataset-specific deep learning models FT-Transformer and ResNet. Even though the AutoML framework AutoGluon shows strong performance values as discussed in Section 5.3, the hyperparameter optimization comes at the cost of higher computational cost resulting in expensive costs w.r.t. the temporal domain. Both TP-Berta and TabPFN are only evaluated on a small subset of the available datasets as indicated by the small cumulative proportion value for both approaches, where the latter shows its strong performance on the datasets it has been evaluated on by a steep increase. The approaches SAINT, TabNet, CARTE, and XTab range in-between, where CARTE and SAINT show slightly better performance-cost ratios compared to AutoGluon with increased values for the performance ratio $\tau$ considering a broader range of various datasets.
The right plot shows the equivalent performance profiles w.r.t. the measured total time. Notably, XGBoost and AutoGluon with hyperparameter optimization result in strong performance-time ratios. Transformer-based models are outperformed by more lightweight models like CatBoost and ResNet which both show competitive results. From the model family encompassing foundation models, XTab shows the strongest performance, however, is outperformed by classical GBDT approaches. TabPFN as an in-context learning model is only applicable on small data regimes and is therefore not competitive considering all datasets. In Appendix E.6, we provide a performance profiles analysis in the small-data and large-data regimes separately.
6 Conclusion
Our comprehensive empirical study evaluates the quality of eleven state-of-the-art tabular classification models. We categorize the approaches into model families according to their underlying learning scheme, which encompasses gradient-boosted decision trees (GBDTs) as tree-based models, foundation models with sub-paradigms of in-context learning and fine-tuning, and dataset-specific neural networks as umbrella for feedforward networks and transformer-based approaches. Our study is conducted with 68 diverse datasets from the OpenMLCC18 benchmark repository. Our study provides a rigorous assessment of state-of-the-art learning paradigms that reveals that dataset-specific neural networks, e.g., ResNet, FT-Transformer, and SAINT, generally outperform meta-learned approaches, e.g., TP-BERTa. We are the first to have a deeper investigation on foundation models, showing that TabPFN excels all other models in small data scenarios. However, classical machine learning methods, such as XGBoost and CatBoost still demonstrate robust performance while highly efficient in comparison to other model families on a broad range of datasets. Moreover, AutoGluon as an AutoML framework exhibits superior performance across diverse datasets, albeit at the cost of increased computational resources. Next to a fair comparison of model families, we provide an in-depth analysis of the influence of hyperparameter optimization on the models’ performance and provide a cost-efficiency analysis that highlights that GBDT approaches outperform most modern deep learning methods. Our study contributes valuable insights into the current landscape of tabular classification data modeling and encourages further potential research directions with promising model families.
Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB 1597.
Appendix A Configuration Spaces
A.1 CatBoost

### v3

@@ 3.2 Experimental Setup
Metrics. Lastly, we report the model’s performance as the average Area Under the Receiver Operating Characteristic (ROC-AUC) across 10 outer test folds. Given the prevalence of imbalanced datasets in the OpenMLCC18 benchmark, we employ ROC-AUC as our primary metric, since it offers a more reliable assessment of model performance.
Code: For reproducibility, our code is available at: https://github.com/machinelearningnuremberg/TabularStudy.
4 Baselines
In our experiments, we compare a range of methods categorized into three distinct groups: Classical Machine Learning Classifiers, Deep Learning Methods, and AutoML frameworks, as shown in Figure 1.
Classical Machine Learning Classifiers. First, we consider XGBoost (Chen & Guestrin, 2016), a well-established GBDT library that uses asymmetric trees. Moreover, we consider CatBoost (Prokhorenkova et al., 2018), a well-known library for GBDT that employs oblivious trees as weak learners and natively handles categorical features with various strategies. Finally, we also include LightGBM (Ke et al., 2017), a widely used GBDT framework that grows trees leaf-wise and supports efficient handling of large datasets.
Deep Learning Methods. In terms of classical deep learning methods, we include the ResNet implementation provided in the work by Gorishniy et al. (2021). Furthermore, we include three recent and competitive variants of the MLP architecture: i) an MLP architecture enhanced with numerical embeddings (Gorishniy et al., 2022) to which we refer as MLP, ii) RealMLP (Holzmüller et al., 2024), an MLP enhanced with several additions like robust scaling, numerical embeddings, etc, iii) TabM (Gorishniy et al., 2025), an efficient ensemble of MLP models.
In terms of transformer-based architectures, we consider TabNet (Arik & Pfister, 2021), an architecture that employs sequential attention to selectively utilize the most pertinent features at every decision step.
Next, we consider SAINT (Somepalli et al., 2021), a hybrid deep learning approach tailored for tabular data challenges. SAINT applies attention mechanisms across both rows and columns and integrates an advanced embedding technique. Lastly, we consider FT-Transformer (Gorishniy et al., 2021), an adaptation of the Transformer architecture for tabular data. It transforms categorical and numerical features into embeddings, which are then processed through a series of Transformer layers.
Foundation Models for Tabular Classification.  For in-context learning, we consider TabPFN (Hollmann et al., 2023), a meta-learned transformer architecture. Next, we consider TabPFNv2 (Hollmann et al., 2025), which alternates attention first across features, then across samples. Finally, we consider TabICL (Qu et al., 2025), pretrained on synthetic datasets similar to the TabPFN models, which can handle up to 500K samples. Among fine-tuned models, we include XTab (Zhu et al., 2023), a method that proposes a cross-table pretraining approach that can work across multiple tables with different column types and structures. Next, we consider TP-BERTa (Yan et al., 2024), a variant of the BERT language model that is adapted for tabular prediction. It introduces a relative magnitude tokenization to transform continuous numerical values into discrete high-dimensional tokens. Lastly, we include CARTE (Kim et al., 2024) in our experimental study. CARTE utilizes a graph representation of tabular data to process tables with differing structures.
Since all the fine-tuned models were pretrained on real-world datasets, we ensure that there are no datasets that overlap with the OpenMLCC18 benchmark.
AutoML Frameworks. Due to the large number of AutoML frameworks available in the community (Feurer et al., 2015; Erickson et al., 2020; LeDell & Poirier, 2020; Feurer et al., 2022), it was infeasible to include all of them in our experimental study. Therefore, we selected AutoGluon (Erickson et al., 2020), a framework that achieves the highest predictive performance in the recent AutoML Benchmark study (Gijsbers et al., 2024).
For all methods, we use their official implementations. We refer the readers to Appendix A for more details.
5 Experiments and Results
Figure 2: Left: Distribution of ranks for the Deep Learning ($12$ methods), Classical ML (33 methods) and AutoML (11 method) classifier families. Right: Distribution of ranks for the Foundation Models (55 methods), Dataset-Specific (77 methods) and AutoML (11 method) classifier families. The boxplots illustrate the rank spread, with medians represented by black lines, diamonds representing the means, and whiskers showing the range.
Research Question 1: Do DL models outperform gradient boosting methods in tabular data classification?
Figure 3: Win-rate dueling matrix comparing learning methods across shared datasets. Each cell (row $i$, column $j$) shows the fraction of common datasets on which method $i$ outperforms method $j$.
To address our research question, we initially compare the performance of Deep Learning methods and Classical ML methods jointly, by ranking the methods per-dataset and analyzing the rank distribution (the lower the rank, the better). The results provided in Figure 2 (Left) indicate that DL methods outperform the previous state-of-the-art GBDTs approaches. The best performing methods are TabICL with a median rank of 2, followed by AutoGluon and TabPFNv2, both with a median rank of 3. In terms of non-meta learned methods, TabM achieves a median rank of 5, followed by CatBoost, XGBoost, and LightGBM, with median ranks of 5.5, 7, and 7.5, respectively.
Next, we investigate how the models compare in a one-versus-one setting to eliminate the effect of related baselines. We present the results in Figure 3, where we observe that the one-versus-one results are consistent with the results where all the methods are considered jointly. In terms of meta-learned architectures, both TabPFNv2 and TabICL outperform tree-based architectures in the majority of the datasets. From the non-meta learned models, only TabM manages to outperform all variants of tree-based architectures.
Figure 4: Dataset landscape showing winning method families across different dataset sizes. Each point represents a dataset from the OpenMLCC18 benchmark, positioned by number of rows (x-axis) and features (y-axis) on log scales. Colors indicate which method family achieved the highest accuracy: Deep Learning methods (orange), Classical ML tree-based models (green), and ties (gray).
Lastly, we investigate whether there exists a certain region where deep learning methods are superior compared to the tree-based baselines, or where the opposite holds. For that purpose, we highlight the winning method family in Figure 4 for every dataset, over the number of examples and number of features of a dataset. The results show that Deep Learning methods dominate tree-based methods in datasets that have less than 5000 examples, by winning 31-3. In cases where a dataset has more than 5000 examples, tree-based methods become more competitive. However, they are still outperformed by deep learning methods 17-7.
For additional analyses, such as evaluations of predictive performance across different data regimes, examinations of the cost–efficiency trade-off, and investigations into the influence of meta-features on predictive performance, we refer readers to Appendices G, H, and I.
Research Question 2: Do meta-learned NNs outperform data-specific NNs in tabular data classification?
To answer the second research question, we analyze the distribution of ranks between the two subfamilies within the Deep Learning category: foundation models and dataset-specific neural networks. Figure 2 (Right) plot illustrates that in-context learning models are very competitive, with TabICL and TabPFNv2 having the best overall rank. Within the dataset-specific family, TabM demonstrates the best performance, attaining a median rank of 4 across all 68 datasets. The fine-tuned foundation models XTab, CARTE, and TP-BERTa obtain the worst performance compared to all remaining dataset-specific neural networks, except for TabNet. Interestingly to note is that, except for the in-context learning models, which are meta-learned architectures, the feed-forward neural networks TabM, RealMLP, and MLP outperform the attention-based architectures.
Next, we compare foundation models with dataset-specific NNs by generating critical difference diagrams. To generate the CD diagrams, we utilize the autorank package (Herbold, 2020), which performs a Friedman test followed by a Nemenyi post-hoc test at a significance level of $0.05$.
Figure 5: Critical difference (CD) diagram of the methods, where a horizontal bar indicates the absence of statistical significance. Left: CD diagram of Deep Learning vs. GBDTs, Right: CD diagram of dataset-specific vs. foundation models.
We present the results in Figure 5, where the black bars connecting methods indicate that there is no statistically significant difference in performance. Due to the limited number of datasets shared among the methods, TP-BERTa was excluded from this comparison. The left diagram of Figure 5 illustrates that TabPFNv2 and TabICL outperform all the other methods, demonstrating superior performance. TabM trails the top methods by a narrow margin, ranking third overall. CatBoost and XGBoost follow, and their performance differences relative to the top three are not statistically significant. We also compare the performance of models within the deep learning family. The right plot of Figure 5 shows a critical difference diagram indicating that TabICL and TabPFNv2 again attain the top average ranks. Except for TabM, both TabICL and TabPFNv2 are significantly better than all remaining methods.
A comprehensive presentation of the raw results for all methods, both after hyperparameter optimization (HPO) and with default hyperparameter configurations, is provided in Appendix D.
Figure 6: Left: Distribution of ranks for the Fine-tuning (3 models) and In-context learning (3 models) classifier families. The boxplots illustrate the rank spread, with medians represented by black lines and whiskers showing the range. Right: CD diagram of Fine-tuning models against In-context learning models. The horizontal bar indicates the absence of statistically significant differences.
Research Question 3: Which paradigm in transfer learning performs better: Do in-context models or fine-tuned models perform better?
To further investigate the family of foundation models, whether fine-tuning or in-context learning models yield better performance, we conducted an analysis similar to our previous research questions. We employ boxplots to display the distribution of ranks and use critical difference (CD) diagrams to evaluate the statistical significance of the results.
Figure 6 (Left) illustrates that TabICL, categorized under in-context learning methods, achieved a median rank of 1, and TabPFNv2 achieved a median rank of 2. Among the fine-tuning methods, XTab showed the best performance with a median rank of 3 but exhibited a larger interquartile range compared to the in-context learning methods, followed by CARTE and TP-BERTa.
To investigate whether the differences in performance are statistically significant, we present a CD diagram in Figure 6 (Right), from which TP-BERTa and TabPFN are excluded due to the limited number of common datasets among the methods. The CD diagram reveals that the in-context learning methods, TabICL and TabPFNv2, significantly outperform the fine-tuning methods XTab and CARTE.
Research Question 4: Does refitting after performing hyperparameter optimization have a significant impact on the predictive quality of the models, and does it impact the overall model ranking?
To investigate the impact of refitting, we select two top-performing distinct methods from the Deep Learning family, namely FT-Transformer as a transformer-based architecture, and TabM as an MLP-based architecture, while also selecting CatBoost and XGBoost from the tree-based family as the top-performing models.
Figure 7: Refitting impact on the predictive performance. Left: Investigating the rank distribution of the methods in isolation, with and without refitting. Right: Investigating the distribution of the ranks for the methods jointly, with and without refitting.
Initially, we compare the models in isolation to investigate how refitting affects the distribution of predictive performances across tasks. We present the results in Figure 7 (Left), where, as observed, all the methods that incorporate refitting feature a lower rank and outperform their non-refitting counterparts. Additionally, as we show in Appendix C, the difference in results is statistically significant in the majority of cases.
Moreover, we investigate whether refitting affects the ranking of the methods when considered jointly. The results in Figure 7 (Right) indicate that refitting does change the ranking of the methods, where, e.g., after refitting, XGBoost manages to outperform FT-Transformer and achieves a better median rank compared to the non-refitting counterpart. We continue by evaluating the impact of refitting via head-to-head, dataset-level comparisons, and applying statistical tests to assess the significance of the observed differences. We kindly refer the reader to Appendix C for details.
Research Question 5: What is the influence of hyperparameter optimization on a method’s predictive performance?
Figure 8: Rank improvement of methods with hyperparameter optimization compared to using the default configuration. Rankings are computed relative to other methods on each dataset (default rank – HPO rank; positive values indicate improvement). Box plots show the distribution across datasets, with points for individual datasets, horizontal bar for the median and diamonds for the mean.
To investigate the impact of hyperparameter optimization, we calculate the per-dataset rank of every method with and without performing hyperparameter optimization and compare the median rank improvement for every method compared to the other baselines. In general, the majority of methods improve in predictive performance when considered in isolation, as validated in Appendix B.13. However, the results in Figure 8 indicate that only TabM, RealMLP, SAINT, XGBoost, and XTab improve in terms of median rank compared to the other contenders, when hyperparameter optimization is performed.
In Appendix B, we provide a detailed analysis of hyperparameter importance, showing both the overall contribution of hyperparameters to model performance and the individual effect of each hyperparameter for every method.
6 Conclusion
Our comprehensive empirical study evaluates the quality of seventeen state-of-the-art tabular classification approaches across 68 diverse classification datasets from the OpenMLCC18 benchmark with a rigorous setup, employing cross-validation, model-based hyperparameter optimization, and refitting.
Our results indicate a paradigm shift, where deep learning methods outperform traditional baselines in all dataset regimes of the considered benchmark. Next to a fair comparison of model families, we provide an in-depth analysis of the importance of refitting, the influence of hyperparameter optimization on the models’ performance, the most important hyperparameters per method, and the cost-performance efficiency of various methods.
Our study contributes valuable insights into the current landscape of tabular data modeling and encourages further potential research directions with promising model families.
Acknowledgments
We acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under SFB 1597 (SmallData), grant number 499552394.
Appendix A Evaluation protocol and Configuration Spaces
A.1 Evaluation Protocol
@@ 22   Train final model $M(\lambda^{*}_{k})$ on $D_{\text{train}}^{k}$;
Algorithm 1, shows the nested-cross validation with the outer folds (lines 1-1) and inner folds (lines 1-1). In each trial (lines 1-1), the mean performance across inner folds are calculated in line 1 which is used as the objective value for Optuna in line 1. After the maximal number of trials $T$ is reached or the time budget is exceeded, we select the best hyperparameter setting in line 1.
A.2 CatBoost